### v1

# Can Automated Feedback Turn Students into Happy Prologians?

Abstract
Giving personalized feedback to students is very important to the learning process. However, doing so in a timely manner can be difficult to accomplish in very large courses. Recent work has explored different types of automated feedback adapted to different languages and programming paradigms, particularly logic programming. In ProHelp, we implemented several of these types of feedback so that they could be used by students enrolled in a logic programming class. Then, we surveyed those students to find if the feedback was useful and which types of feedback they preferred. Results show that students found all types of feedback helpful, with automatic testing, in particular, being the most helpful type. We also explore student preferences for which types of feedback they would most like to see implemented in the future.
Keywords:
Automated Feedback Logic Programming User Survey Computer-aided Education.
1 Introduction
Giving timely and formative feedback to students is very important for the learning process [11, 28]. However, doing so in large courses is a difficult task [19], and even impossible in cases like Massive Open Online Courses (MOOCs). Over the years, many automated assessment frameworks have been proposed to address this problem. Automated assessment tools such as Mooshak/Enki [20] or GitSEED [18], among others [6, 10, 22, 26], allow faculty to automatically test students’ submissions and provide them with varying levels of feedback.
Many newer automated assessment systems focus on (1) making it more usable for students by targeting well-known interfaces such as Git, and (2) making it easier to maintain and modify by faculty. Such tools, like GitSEED, allow faculty to fully customize what feedback is returned to the users. This makes it easy to exploit state-of-the-art software analysis and automated diagnosis tools, which in turn allow students to more quickly find their mistakes and repair them.
Novice students usually find it hard to debug logic programs. This is mainly due to the lack of traditional control and data flows they learn in imperative languages. Hence, we propose ProHelp, an automated feedback tool for Prolog that can be integrated with GitSEED. ProHelp includes feedback features that range from automatic testing to typo correction suggestions and validation of implementation techniques.
The two main goals of this work are to (1) understand which types of feedback implemented in ProHelp students find the most helpful, and (2) gauge students’ interest in more complex types of automated feedback they would like to see in the future. To this end, we designed and deployed a survey with the students that interacted with ProHelp. Results show that the feedback provided to students was useful and the system has good usability. Students also showed great interest in proposed improvements and suggested some themselves.
2 Related Work
2.0.1 Automated Assessment Tools (AATs).
Automated assessment of programming assignments has a long history of more than 60 years of research [19]. Most Automated Assessment Tools assess programs by either comparing them with a known-correct solution or by executing a series of tests and checking their results [11]. A large number of current AATs are based around Integrated Development Environment-like web-interfaces with examples like CodeOcean [26], Mooshak/Enki [14, 20] and Web-Cat [6]. Other platforms, like ProgEdu [3], GitHub Classroom [^1] https://classroom.github.com/ and GitSEED [18] focus on Git, taking advantage of Continuous Integration (CI) actions to assess students’ work as it is submitted to the repository. These tools can be easier for students to use because they are often already familiar with the Git workflow.
2.0.2 Automated Feedback for Logic Programming.
Many systems have been proposed for automated feedback, debugging and repair of logic programming languages [13, 16, 27, 1]. These tools face different challenges compared with other programming paradigms due to the declarative nature of the languages, lack of explicit control flow and issues like non-determinism, non-monotonic reasoning or non-termination [16]. INCOM [13] introduces a two-stage tutoring model where students first engage in task analysis progressing to the implementation of Prolog predicates. This structured approach assists learners in bridging the gap between problem specification and algorithm development. Meanwhile, PRAM [16] focuses on the automatic assessment of Prolog programs by incorporating both static and dynamic metrics to evaluate code style, complexity, and dynamic correctness, thereby reducing instructor workload and facilitating the shift from procedural to declarative thinking. Complementing these educational tools, ProFL [27] applies fault localization techniques, particularly coverage-based and mutation-based approaches, to identify errors in Prolog programs effectively. Moreover, the FormHe [1] system extends automated feedback to Answer Set Programming by combining traditional fault localization methods with modern strategies like Large Language Model (LLM)-based fault localization and repair, and mutation-based program repair techniques.
One common element of earlier works on automated logic programming feedback is the use of structured, rule based feedback, which can force students to solve problems in a certain way. Meanwhile, more recent approaches follow a testing-based fault localization method which helps students find problems in their programs instead of guiding them in a particular direction. However, these newer tools lack comprehensive user studies to evaluate their performance. Our work addresses this research gap by directly engaging with the target audience of the tool – the students – and prioritizing their needs.
3 Automated Assessment System
At the core of this work is ProHelp, a new tool for automatic feedback to students learning logic programming with Prolog. ProHelp was integrated with GitSEED [18], a recent easy-to-use automated assessment tool based on GitLab. In this section, we present the overall framework, as well as the feedback extensions we implemented on top of GitSEED.
3.1 GitSEED
Figure 1: Simplified overview of GitSEED’s operation scheme.
GitSEED [18] is an open-source, language-agnostic automated assessment tool integrated with GitLab. It provides students with personalized feedback on programming assignments while they learn Git fundamentals. Unlike traditional assessment platforms, GitSEED operates within GitLab’s CI workflow, eliminating the need for students to learn new interfaces.
As Figure 1 shows, when students submit their work to GitLab, GitSEED automatically evaluates it using a predefined test suite and program analysis tools specified by the faculty. The evaluation report is then pushed directly to the students’ git repositories, ensuring immediate access to personalized feedback.
3.2 Automatic feedback for Prolog submissions
(a) Automatic testing
(b) Predicate Scoring
(c) Syntax error highlighting
(d) Open choice points
(e) Score rankings
(f) Solution type validation
(g) Unknown predicate name suggestions
Figure 2: Screenshots of the different types of student feedback available in ProHelp.
The minimum configuration of GitSEED simply runs a test suite on the student’s submission and reports the results as feedback. However, as shown in Figure 1, GitSEED can be extended with custom program analysis tools that give students more personalized feedback. Taking advantage of this customizability, we implemented several types of feedback in ProHelp that are specific to the Prolog context. Below, we describe each of the 7 types of automated feedback, along with an example for each one, shown in Figure 2.
  1. a)
    Automatic testing:  This is the basic feature consisting of running the test suite and reporting the results. The faculty can decide to show the body of failing test cases or not.
  2. b)
    Predicate scoring:  For graded assignments, each predicate $p$ that the students need to implement gets assigned X points by the faculty. If a student passes the tests for $p$ then they get X points.
  3. c)
    Syntax error highlighting:ProHelp highlights snippets of code with syntax errors, which makes them easier to locate compared with the line and column coordinates provided by the Prolog interpreter.
  4. d)
    Open choice points warning:  Prolog allows enumerating multiple answers for a given query. However, students often forget this fact. This can result in the first answer being correct w.r.t. the specification but one of the following answers being incorrect. When a predicate terminates with an open choice point (indicating that there are more possible answers) we give a warning to students so that they can validate if this is intended behavior.
  5. e)
    Score rankings:  Each assignment has a global leaderboard that shows the scores and number of passed tests for each student. This incentivizes some students to compete and solve the most exercises as early as possible.
  6. f)
    Solution type validation:  Some of the assignments ask students to solve problems using specific techniques (for example, using recursion). In these assignments, we use a heuristic to determine if the student’s program is using the correct method or not.
  7. g)
    Unknown predicate name suggestions:  When a student uses a predicate that is undefined, ProHelp suggests other predicates with similar names if they exist.
3.3 Future directions
One of the main goals of this work is to gauge students’ opinions on the feedback features already implemented to decide which ones to prioritize for future work. Planned features include several AI-based procedures using formal methods targeted to the specific context of logic programming.
  • Fault localization:  Identifying what section of the program contains bugs could help students narrow their focus. The granularity of the bug location could also be controlled by faculty such as identifying a specific predicate, or a specific set of lines.
  • Bug fix suggestions:  The bug identification can be taken one step further and also include a description of the type of bug and/or a hint on how to fix it. One way to accomplish this is to use automated program repair to find a correction of the bug and then combine the buggy program and the corrected program to generate a hint.
  • Infinite loop identification:  Prolog programs can suffer from infinite loops due to the depth-first search algorithm used by default in many Prolog interpreters. Helping students identify which sections of the program are responsible for the infinite loop behavior can be a helpful tool for students.
  • Test output differences:  The test suites in ProHelp are composed of unit tests implemented through plunit. Almost always, these tests consist of a series of predicate calls followed by an assertion that compares a value with the expected response. It would be useful for students to provide more information when these assertions fail, particularly showing the difference between generated and expected values.
4 Methodology
ProHelp was deployed in a BSc-level class on logic reasoning and programming with 365 students at University X [^2] Anonymized for double-blind review.. During the course, students were given 3 different types of Prolog assignments: a 5-part story-driven introductory exercise, recitation exercises and a course project. Of these, the first two types were optional and ungraded, while the project was mandatory and represented 50% of the final grade of the course.
Overall, 312 students interacted with ProHelp during the course. We obtained 148 responses to our survey (47%), of which we dropped 4 due to failed attention checks [17]. Of the 144 valid responses, 28 respondents identified as female (19%), 111 as male (77%), 1 as other (<1%) and 4 preferred not to disclose (3%). While we have no information about the gender distribution for students who took this class, the official distribution for the enrolled students in this BSc degree is 83% male students and 17% female. This provides some confidence that our sample is representative of the overall course population.
To understand the strengths and shortcomings of ProHelp, we performed a post-experience survey. The survey had an estimated duration of 15 minutes and was performed using EUSurvey [^3] https://ec.europa.eu/eusurvey, during the two weeks following the last programming assignment. With this survey, we plan to answer the following three research questions:
  1. RQ1.
    How valuable did students find the feedback, and which aspects of it were the most useful?
  2. RQ2.
    Does student interest, engagement or LLM-usage affect how helpful they found the feedback?
  3. RQ3.
    What features do students believe would be most beneficial in the future?
4.1 Data Analysis
Our study focuses on analyzing 3- and 5-point Likert scales regarding the usefulness of different aspects of the feedback received. For RQ1 and RQ3, we are interested in how different answers of the same student compare to one another (the ratings for the different types of feedback). To accomplish this, we use the Friedman rank sum test [7], to check for any statistically significant differences between the feedback types. In cases where there are such differences, we employ the post-hoc pairwise Durbin-Conover test [4] with the Holm correction method [9], in order to identify which pairs of feedback types have statistically significant differences in usefulness.
In RQ2, we are interested in how different sub-groups of students rated the usefulness of ProHelp (on a 5-point Likert scale). To accomplish this, we use different tests depending on the nature of the independent variable. Students’ interest was measured on a scale from 1 to 5 (an ordinal variable) and, as such, we use the Spearman’s rank correlation [25] to compare it with the usefulness. Meanwhile, students’ engagement and usage of LLMs were both measured through categorical variables. We use the Kruskal-Wallis test [8] for the engagement since it has more than 2 categories, and the Mann-Whitney $U$ test [15] for the LLM-usage since it only has 2. The significance level for all tests was set at $0.05$.
Finally, our survey also contained two open-ended questions regarding future feedback suggestions and general opinions on the system. The answers were coded according to a unified code-book for Feedback and Improvements. The code-book was created by author X. The answers were coded, independently, by authors X and Y, with any differences conciliated through a posterior discussion.
4.2 Ethics and Privacy
All students had access to the same system with the same types of feedback available. This ensured that no group of students had any advantage over the rest. Participation in this study was voluntary, unremunerated, and all participants provided informed consent for data collection and processing before the study. Participants were recruited through a course announcement after the end of the last assignment. No identifiable personal data was collected. In University X, where the course was located and the survey performed, studies with these characteristics do not need to be submitted to the Ethics Committee (as confirmed by the Chair of the Ethics Committee). Co-authors from other institutions than University X only had access to aggregate statistics and not individual responses. We have completed the European Commission’s ethics self-assessment checklists and comply with all the guidelines.
5 Results
This work aims to understand the advantages and limitations of the ProHelp system and explore what types of feedback students would like to have in the future. The statistical analysis and plots presented in this section were created using R version 4.4.2 [23], statsExpressions [21], ggstats [12] and ggsignif [5].
Figure 3: Agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
5.1 Usefulness and usability of the feedback (RQ1)
To gauge if ProHelp was useful to students, we asked them if they agreed with the statement “I think the automated feedback helped increase my grade in this course” using a 5-point Likert scale. Figure 3 shows that 74% of student responses agreed that the provided feedback was useful, while 11% disagreed and 15% neither agreed nor disagreed. This indicates that ProHelp was helpful for students. We also performed a usability analysis through the System Usability Scale (SUS) questionnaire [^4] We used a peer reviewed translation of the SUS questions. The language and reference have been omitted for anonymity and will be included in the final version. [2]. The average SUS score among our respondents was 78.5 which corresponds to a B+ grade according to the curved grading interpretation by Sauro and Lewis [24]. While there is room for improvement, the results show that our system is fairly usable in its current state.
Figure 4: Usefulness rating of the different types of feedback. The median response for each type is shown by the vertical line. The right side of the figure shows pairs of feedback types for which the Durbin-Conover post-hoc test indicates statistically significant differences in ratings. For example, there is a statistically significant difference between the responses for “Automatic testing” and “Score rankings” with $p=1.23\mathrm{e}\text{-}07$.
This study also aims to understand if any type of feedback stood out as particularly more useful than the others. Hence, we asked the students to rank each of the 7 types of feedback from Section 3.2 from 1 (not useful) to 5 (very useful). Figure 4 presents the answer distribution to these questions. Most students ranked all 7 types of feedback as useful, with all of the medians being $\geq 4$.
Since not all students interacted with all forms of feedback and the Friedman test requires a full block design, we excluded students who did not answer all 7 questions, resulting in 94 samples. The Friedman test allows us to reject the null hypothesis of there being no significant difference between the different types of feedback ($\chi^{2}(6)=58.66,\;p=8.436\times 10^{-11},\;\hat{W}_{Kendall}=0.10$). As such, to discover which types of feedback were significantly more useful, we used the post-hoc pairwise Durbin-Conover test. The right side of Figure 4 shows the pairs of feedback types for which there is a significant difference in median ranks, along with the respective p-values. These statistically significant tests show that “Automatic testing” was the most helpful type of feedback, being more useful than 4 other types, and that “Open choice points warning” and “Predicate scoring” were more useful than 2 other types each. Meanwhile, the least helpful types of feedback were “Score rankings” and “Solution type validation”, with each being less useful than 3 other types.
Most students consider that the provided feedback was helpful and the system has a usability score of B+. The most useful feedback were “Automatic testing”, “Open choice points warning” and “Predicate scoring”, while the least useful were “Score ranking” and “Solution type validation”. These conclusions are supported by statistically significant differences.
5.2 Impact of interest, engagement and LLM use (RQ2)
Another goal of this study is to determine if students’ interest, engagement with optional exercises and having used LLMs (or not) had any effect on their perception of the usefulness of the feedback.
5.2.1 Interest
Figure 5: Answer distribution for the question “How interested are you in the topics of logic and logic programming?”. The median response is shown by the black vertical line.
We asked students to rate their interest in the topics of logic and logic programming on a scale from 1 to 5. The answers, summarized in Figure 5, show a fairly normal distribution, with most students having some interest in the topics and a small number showing no interest or a large amount of interest.
Figure 6: Effect of interest in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
To test our hypothesis that more interested students take better advantage of automated feedback, we decided to compare students’ interest with the agreement level for the statement “I think the automated feedback helped increase my grade in this course”. Figure 6 shows the agreement responses grouped by interest level. A Spearman’s rank correlation analysis does not allow us to reject the null hypothesis, which means there is no statistically significant relationship between students’ interest level and how beneficial they found the feedback received ($S=4.26\times 10^{5},\;p=0.08,\;\hat{\rho}=0.14$).
5.2.2 Engagement
Figure 7: Effect of engagement in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
Students had the opportunity to use ProHelp with two types of optional exercises: a set of 5 introductory exercises and a set of 22 recitation exercises. Considering the 144 respondents to our survey, 47 did not solve any optional exercises, 50 solved at least one exercise from both optional sets, 21 solved only exercises from the introductory set, and 26 solved only exercises from the recitation set. Our hypothesis is that more engaged students (ones who solved optional exercises) would have taken better advantage of ProHelp and thus found it more useful. Figure 7 shows the relationship between the 4 engagement groups and the agreement levels with the statement “I think the automated feedback helped increase my grade in this course”. To test our hypothesis, we used a Kruskal-Wallis test, which shows no evidence of there being a statistical difference between the four groups ($\chi^{2}(3)=2.19,\;p=0.53,\;\hat{\epsilon}^{2}=0.02$).
5.2.3 Usage of LLMs
Figure 8: Effect of LLMs usage in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
Students were allowed to use Large Language Models (LLMs) during this course, as long as they disclosed that usage. Considering the 144 responses to our survey, 53 reported using LLMs when solving assignments, while 91 reported that they did not use LLMs. We were interested to know if having used LLMs has an effect on how students perceived the usefulness of ProHelp (for example, students who used LLMs might have found ProHelp less useful due to already having personalized feedback from the LLM). In Figure 8, we show the usefulness ratings for students who used and did not use LLMs. According to the Mann-Whitney $U$ test, there is no statistical difference between the two groups ($W=2261.50,\;p=0.51,\;r=-0.06$).
There is no evidence that interest level, having used the platform for optional exercises, or having used LLMs, affects how helpful students found ProHelp overall.
5.3 Future direction (RQ3)
Figure 9: Usefulness ratings for the different types of feedback not yet implemented in ProHelp. The median response for each type of feedback is shown by the black vertical line. The right side of the figure shows pairs of feedback types for which the Durbin-Conover post-hoc test indicates statistically significant differences in ratings.
Our last goal is to better understand what features and improvements students would most like to have in the future, while also analyzing what aspects of the course students did not like. In the survey, we asked students to rate the 4 types of (not yet implemented) feedback described in Section 3.3 between Not Useful, Somewhat Useful and Useful. Figure 9 shows the students’ answers, where at least 75% of students rated each of the proposed feedback types as Useful. These responses hint that these are good goals to work towards.
To check if any type of feedback had a significant preference over another, we used a Friedman test which rejects the null hypothesis ($\chi^{2}(3)=20.74,\;p=1.19\times 10^{-4},\;\hat{W}_{Kendall}=0.05$). Next, to check which types of feedback students think will be more useful, we used the post-hoc pairwise Durbin-Conover test. The right side of Figure 9 shows the pairs of feedback types that showed statistical significance. Students rated “Showing the differences between generated and expected output” as significantly more useful than the other 3 types of feedback.
At the end of the survey, we asked students to suggest other types of feedback they would like to see in the future, as well as provide general comments/criticisms. One common criticism among students was that the unit tests were not specific enough, i.e., each test covered too much functionality. This left students with little information when tests failed and highlights the importance of faculty creating good unit tests that assess each functionality in isolation. There were also other comments regarding the course organization and grading method, which we have relayed to the faculty.
One interesting suggestion provided by some students regards the qualitative component of the project grade, which is manually attributed and encompasses characteristics such as code readability, organization and documentation. Students suggest that it would be useful to have an estimate for this qualitative component in the automated feedback, even if the final score continued to be hand-assigned by faculty.
Although students found all proposed types of feedback useful, there is a significant preference for the “Show the outputs and expected outputs” compared with the remaining feedback types. Students would like to have an automated estimation for the qualitative component of their projects.
6 Discussion and Conclusion
This work proposes ProHelp, a new tool that provides automated feedback for students learning logic programming using Prolog. Additionally, we explore more advanced types of automated feedback that have not yet been implemented. Our main contribution is a student survey with 144 valid responses, gathering insights about students’ preferences and complaints regarding automated assessment tools, specifically GitSEED and ProHelp.
Survey analysis indicates that users found the feedback received useful and the GitSEED interface was user-friendly. The components of ProHelp that students most appreciated include automatic testing, warnings about open choice points, and scoring for each required predicate. Students also displayed great interest in the more advanced types of feedback we proposed and suggested some themselves. These findings will be valuable for future research, enabling educators and developers to focus on the most impactful types of feedback.
Results show no statistically significant relationship between students’ perception of feedback usefulness and their interest, engagement, or usage of LLMs. Even so, there is a small but non-statistically significant correlation ($p=0.08$) between students’ interest and their usefulness rating, suggesting that further research – with a larger sample or different methodology – is warranted. The findings on LLM usage are particularly intriguing, and we propose several possible explanations for it: (1) students under-reported LLM-usage due to its negative connotation, (2) the feedback received in ProHelp is orthogonal to the questions students usually ask LLMs, or (3) LLMs are not (yet) very useful for Prolog. Further investigation is needed to determine which, if any, of these factors apply.
Building on the results of the survey, we plan to enhance ProHelp with the following improvements: (1) use fault localization and program diagnosis techniques to give more information to students when tests fail, particularly for failed assertions and infinite loops, (2) explore automated repair methods, namely program synthesis and LLMs, to generate correction hints or bug explanations for students, and (3) investigate LLM-based techniques to estimate qualitative aspects of logic programs. One challenge highlighted by some students is that excessive feedback may hinder learning rather than help. Balancing the amount and type of feedback will be crucial to ensuring that students receive meaningful guidance without becoming overly reliant on automation.
In conclusion, this work introduces ProHelp, a tool offering basic automated feedback for Prolog, and presents findings from a user study on students’ preferences for automated assessment tools. The results indicate that students found ProHelp valuable, particularly its automatic testing and reporting features, and showed strong interest in future enhancements.