### v2

# Can Automated Feedback Turn Students into Happy Prologians?

Abstract.
Providing valuable and personalized feedback is essential for effective learning, but delivering it promptly can be challenging in large-scale courses. Recent research has explored automated feedback mechanisms across various programming languages and paradigms, including logic programming.
In this work, we present a student survey were we evaluate the perceived usefulness of different feedback types and identified which are most valued. Our results indicate that students found all implemented feedback types helpful, with automatic testing ranked as the most useful. We also introduce a dataset comprising 7201 correct and incorrect Prolog submissions, along with 200 manually annotated programs labeled with bug types and corresponding corrections. Finally, we explore student preferences for which types of feedback they would most like to see implemented in the future.
Automated Feedback, Logic Programming, User Survey, Dataset, Computer-aided Education
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Applied computing Computer-assisted instructionccs: Human-centered computing Empirical studies in HCIccs: Theory of computation Constraint and logic programming
1. Introduction
Giving timely and formative feedback to students is very important for the learning process (Keuning et al., 2019; Wisniewski et al., 2020). However, doing so in large courses is a difficult task (Paiva et al., 2022), and even impossible in cases like Massive Open Online Courses (MOOCs). Over the years, many automated assessment frameworks have been proposed to address this problem. Automated assessment tools such as Mooshak/Enki (Paiva et al., 2016) or GitSEED (Orvalho et al., 2024), among others (Edwards and Pérez-Quiñones, 2008; Iddon et al., 2023; Peveler et al., 2017; Staubitz et al., 2016), allow faculty to automatically test students’ submissions and provide them with varying levels of feedback.
Many newer automated assessment systems focus on (1) making it more usable for students by targeting well-known interfaces such as Git, and (2) making it easier to maintain and modify by faculty. Such tools, like GitSEED, allow faculty to fully customize what feedback is returned to the users. This makes it easy to exploit state-of-the-art software analysis and automated diagnosis tools, which in turn allow students to more quickly find their mistakes and repair them.
However, despite these advances, significant challenges remain in the context of logic programming. Novice students usually find it hard to debug logic programs. This is mainly due to the lack of traditional control and data flows they learn in imperative languages. In this work, we aim to better understand these difficulties and students’ behavior when learning Prolog. To accomplish this, we created ProHelp, an automated assessment platform for Prolog built on top of GitSEED. ProHelp includes feedback features that range from automatic testing to typo correction suggestions and validation of implementation techniques.
We designed and deployed a mixed-methods study with the students that interacted with ProHelp. We used survey responses and submission logs collected during a 9-week undergraduate course with 365 students. Overall, we analyzed 7201 submissions and 144 survey responses.
The main contributions of this work are:
  • Understanding which types of feedback implemented in ProHelp students find the most helpful;
  • Understanding what types of bugs students have in their submissions;
  • A dataset of 7201 correct and incorrect Prolog programs, with 200 programs manually annotated with bug types and corrections;
  • Gauging students’ interest in more complex types of automated feedback they would like to see in the future.
Results show that the feedback provided to students was helpful, and that the automated assessment system has good usability. Students also showed great interest in proposed improvements and suggested some themselves.
This paper is organized as follows: in Section 2 we present an overview of related work. Following that, in Section 3, we present the course during which this study was performed and in Section 4 describe the automated assessment system and the types of feedback provided. Then, in Section 5, we present the methodology for our study and in Section 6, we present the analysis and results. Finally, in Section 7, we present a discussion on the obtained results and conclude the paper.
2. Related Work
Automated Assessment Tools (AATs).
Automated assessment of programming assignments has a long history of more than 60 years of research (Paiva et al., 2022). Most Automated Assessment Tools assess programs by either comparing them with a known-correct solution or by executing a series of tests and checking their results (Keuning et al., 2019). A large number of current AATs are based around Integrated Development Environment-like web-interfaces with examples like CodeOcean (Staubitz et al., 2016), Mooshak/Enki (Leal and Silva, 2003; Paiva et al., 2016) and Web-Cat (Edwards and Pérez-Quiñones, 2008). Other platforms, like ProgEdu (Chen et al., 2017), GitHub Classroom [^1] https://classroom.github.com/ and GitSEED (Orvalho et al., 2024) focus on Git, taking advantage of Continuous Integration (CI) actions to assess students’ work as it is submitted to the repository. These tools can be easier for students to use because they are often already familiar with the Git workflow.
Automated Feedback for Logic Programming.
Many systems have been proposed for automated feedback, debugging and repair of logic programming languages (Le and Pinkwart, 2011; Mansouri et al., 1998; Thompson and Sullivan, 2020; Brancas et al., 2025a). These tools face different challenges compared with other programming paradigms due to the declarative nature of the languages, lack of explicit control flow and issues like non-determinism, non-monotonic reasoning or non-termination (Mansouri et al., 1998). INCOM (Le and Pinkwart, 2011) introduces a two-stage tutoring model where students first engage in task analysis progressing to the implementation of Prolog predicates. This structured approach assists learners in bridging the gap between problem specification and algorithm development. Meanwhile, PRAM (Mansouri et al., 1998) focuses on the automatic assessment of Prolog programs by incorporating both static and dynamic metrics to evaluate code style, complexity, and dynamic correctness, thereby reducing instructor workload and facilitating the shift from procedural to declarative thinking. Complementing these educational tools, ProFL (Thompson and Sullivan, 2020) applies fault localization techniques, particularly coverage-based and mutation-based approaches, to identify errors in Prolog programs effectively. Moreover, the FormHe (Brancas et al., 2025a) system extends automated feedback to Answer Set Programming by combining traditional fault localization methods with modern strategies like Large Language Model (LLM)-based fault localization and repair, and mutation-based program repair techniques.
One common element of earlier works on automated logic programming feedback is the use of structured, rule-based feedback, which can force students to solve problems in a certain way. Meanwhile, more recent approaches follow a testing-based fault localization method which helps students find problems in their programs instead of guiding them in a particular direction. However, these newer tools lack comprehensive user studies to evaluate their performance. Our work addresses this research gap by directly engaging with the target audience of the tool – the students – and prioritizing their needs.
3. Course Structure
We evaluated student’s behavior during a 9 week-long bachelor’s level logic programming course. During this course, students learned classic logic reasoning, as well as logic programming through Prolog. Overall, students faced three different types of Prolog exercises during this course: a 5-part role-playing exercise to familiarize students with the language, 22 optional recitation exercises, and one graded and mandatory final project. The mandatory project was worth 50% of the final grade, while the remaining 50% was from other theoretical assessments. Next, we describe the Prolog exercises in more detail.
  • Role-playing Exercise
    • Optional, not graded, online only
    • 5 puzzles over 5 days
    • A series of simple logic puzzles where students help a detective catch a group of criminals using logic programming;
  • Recitation Exercises
    • Optional, not graded, partially solved in-class using pen and paper
    • 22 exercises
    • 4 Prolog exercise sheets grouped into the topics of: lists, arithmetic, negation and higher-order functions;
  • Project
    • Mandatory, graded
    • 5 weeks duration
    • Worth 50% of the final course grade. The goal of this year’s project was to create a solver for the logic puzzle star battle;
[…]
Figure 1. Course exercises timeline.
In Figure 1, we show the timeline of the different exercises solved by the students. During the first weeks, the students learned logic fundamentals. Then, during week 3, the students started learning Prolog, with the optional exercise sets being released during weeks 3, 4 and 5. Meanwhile, the project statement was released at the beginning of week 4 and lasted until the end of week 9.
4. ProHelp: Prolog Automated Assessment Platform
Figure 2. Simplified overview of GitSEED’s operation scheme.
The submission and evaluation platform used in this course was ProHelp, an extension built on top of the open-source automated assessment framework GitSEED (Orvalho et al., 2024). GitSEED provides students with personalized feedback on programming assignments while they learn Git fundamentals. Unlike traditional assessment platforms, GitSEED operates within GitLab’s CI workflow, eliminating the need for students to learn new interfaces. As Figure 2 shows, when students submit their work to GitLab, GitSEED automatically evaluates it using a predefined test suite and program analysis tools specified by the faculty. The evaluation report is then pushed directly to the students’ git repositories, ensuring immediate access to personalized feedback.
(a) Automatic testing
(b) Predicate Scoring
(c) Syntax error highlighting
(d) Open choice points
(e) Score rankings
(f) Solution type validation
(g) Unknown predicate name suggestions
Figure 3. Screenshots of the different types of student feedback available in ProHelp.
4.1. Automated Feedback
The minimum configuration of GitSEED simply runs a test suite on the student’s submission and reports the results as feedback. However, as shown in Figure 2, GitSEED can be extended with custom program analysis tools that give students more personalized feedback. Taking advantage of this customizability, we created ProHelp and implemented seven types of feedback that are specific to the Prolog context. Below, we describe each of the 7 types of automated feedback, along with an example for each one, shown in Figure 3.
  1. a)
    Automatic testing:  This is the basic feature consisting of running the test suite and reporting the results. The faculty can decide to show the body of failing test cases or not.
  2. b)
    Predicate scoring:  For graded assignments, each predicate $p$ that the students need to implement gets assigned X points by the faculty. If a student passes the tests for $p$ then they get X points.
  3. c)
    Syntax error highlighting:ProHelp highlights snippets of code with syntax errors, which makes them easier to locate compared with the line and column coordinates provided by the Prolog interpreter.
  4. d)
    Open choice points warning:  Prolog allows enumerating multiple answers for a given query. However, students often forget this fact. This can result in the first answer being correct w.r.t. the specification but one of the following answers being incorrect. When a predicate terminates with an open choice point (indicating that there are more possible answers) we give a warning to students so that they can validate if this is intended behavior.
  5. e)
    Score rankings:  Each assignment has a global leaderboard that shows the scores and number of passed tests for each student. This incentivizes some students to compete and solve the most exercises as early as possible.
  6. f)
    Solution type validation:  Some of the assignments ask students to solve problems using specific techniques (for example, using recursion). In these assignments, we use a heuristic to determine if the student’s program is using the correct method or not.
  7. g)
    Unknown predicate name suggestions:  When a student uses a predicate that is undefined, ProHelp suggests other predicates with similar names if they exist.
4.2. Future Directions
One of the goals of this work is to gauge students’ opinions on the feedback features already implemented and decide which ones to prioritize for future work. Planned features include several AI-based procedures using formal methods targeted to the specific context of logic programming.
  • Fault localization:  Identifying what section of the program contains bugs could help students narrow their focus. The granularity of the bug location could also be controlled by faculty such as identifying a specific predicate, or a specific set of lines.
  • Bug fix suggestions:  The bug identification can be taken one step further and also include a description of the type of bug and/or a hint on how to fix it. One way to accomplish this is to use automated program repair to find a correction of the bug and then combine the buggy program and the corrected program to generate a hint.
  • Infinite loop identification:  Prolog programs can suffer from infinite loops due to the depth-first search algorithm used by default in many Prolog interpreters. Helping students identify which sections of the program are responsible for the infinite loop behavior can be a helpful tool for students.
  • Test output differences:  The test suites in ProHelp are composed of unit tests implemented through plunit. Almost always, these tests consist of a series of predicate calls followed by an assertion that compares a value with the expected response. It could be useful for students to provide more information when these assertions fail, particularly showing the difference between generated and expected values.
5. Methodology
ProHelp was deployed in a BSc-level class on logic reasoning and programming with 365 students at Universidade de Lisboa. During the course, students were given automated feedback on 3 different types of Prolog assignments: a role-playing exercise, recitation exercises and a final project.
To understand the strengths and shortcomings of ProHelp, we performed a post-experience survey. The survey, which is replicated in Appendix A, had an estimated duration of 15 minutes and was performed using EUSurvey [^2] https://ec.europa.eu/eusurvey during the two weeks following the last programming assignment.
Of the 312 students that interacted with ProHelp during the course, we obtained 148 responses to our survey (47%), of which we dropped 4 due to failed attention checks (Oppenheimer et al., 2009). Of the 144 valid responses, 28 respondents identified as female (19%), 111 as male (77%), 1 as other (¡1%) and 4 preferred not to disclose (3%). While we have no information about the gender distribution for students who took this class, the official distribution for the enrolled students in this BSc degree is 83% male students and 17% female. This provides some confidence that our sample is representative of the overall course population.
In this work, we plan to answer the following four research questions:
  1. RQ1.
    How valuable did students find the feedback, and which aspects of it were the most useful?
  2. RQ2.
    Does student interest, engagement or LLM-usage affect how helpful they found the feedback?
  3. RQ3.
    What types of bugs did students introduce in their Prolog programs?
  4. RQ4.
    What features do students believe would be most beneficial in the future?
5.1. Data Analysis
Part of our study focuses on analyzing 3- and 5-point Likert scales regarding the usefulness of different aspects of the feedback received. For RQ1 and RQ4, we are interested in how different answers of the same student compare to one another (the ratings for the different types of feedback). To accomplish this, we use the Friedman rank sum test (Friedman, 1937), to check for any statistically significant differences between the feedback types. In cases where there are such differences, we employ the post-hoc pairwise Durbin-Conover test (Conover, 1999) with the Holm correction method (Holm, 1979), in order to identify which pairs of feedback types have statistically significant differences in usefulness.
In RQ2, we are interested in how different sub-groups of students rated the usefulness of ProHelp (on a 5-point Likert scale). To accomplish this, we use different tests depending on the nature of the independent variable. Students’ interest was measured on a scale from 1 to 5 (an ordinal variable) and, as such, we use the Spearman’s rank correlation (Spearman, 1904) to compare it with the usefulness. Meanwhile, students’ engagement and usage of LLMs were both measured through categorical variables. We use the Kruskal-Wallis test (Hollander et al., 2013) for the engagement since it has more than 2 categories, and the Mann-Whitney $U$ test (Mann and Whitney, 1947) for the LLM-usage since it only has 2. The significance level for all tests was set at $0.05$.
The survey also contained two open-ended questions regarding future feedback suggestions and general opinions on the system. The answers were coded according to a unified code-book for Feedback and Improvements. The code-book was created by the first author. The answers were coded, independently, by the first and second authors, with any differences conciliated through a posterior discussion.
At the end of the course, students were given the option to consent to the use of their code for research purposes. The submissions from students who provided consent were fully anonymized prior to any analysis or code inspection. A subset of these submissions was manually examined to categorize the types of bugs students introduced. The buggy programs were labeled using a hierarchical, emergent coding scheme created by the first author.
5.2. Ethics and Privacy
All students had access to the same system with the same types of feedback available. This ensured that no group of students had any advantage over the rest. Participation in the survey was voluntary, unremunerated, and all participants provided informed consent for data collection and processing before the survey. Participants were recruited through a course announcement after the end of the last assignment. No identifiable personal data was collected. In Universidade de Lisboa, where the course was located and the survey performed, studies with these characteristics do not need to be submitted to the Ethics Committee (as confirmed by the Chair of the Ethics Committee). Co-authors from other institutions than Universidade de Lisboa only had access to aggregate statistics and not individual responses. We have completed the European Commission’s ethics self-assessment checklists and comply with all the guidelines.
6. Results
This work aims to understand the advantages and limitations of the ProHelp system and explore what types of feedback students would like to have in the future. The statistical analysis and plots presented in this section were created using R version 4.4.2 (R Core Team, 2024), statsExpressions (Patil, 2021), ggstats (Larmarange, 2025) and ggsignif (Constantin and Patil, 2021).
Figure 4. Agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
6.1. Usefulness and usability of the feedback (RQ1)
To gauge if ProHelp was useful to students, we asked them if they agreed with the statement “I think the automated feedback helped increase my grade in this course” using a 5-point Likert scale. Figure 4 shows that 74% of student responses agreed that the provided feedback was useful, while 11% disagreed and 15% neither agreed nor disagreed. This indicates that ProHelp was helpful for students. We also performed a usability analysis through the System Usability Scale (SUS) questionnaire [^3] We used a peer reviewed translation of the SUS questions. The language and reference have been omitted for anonymity and will be included in the final version. (Brooke, 1996). The average SUS score among our respondents was 78.5 which corresponds to a B+ grade according to the curved grading interpretation by Sauro and Lewis (Sauro and Lewis, 2016). While there is room for improvement, the results show that our system is fairly usable in its current state.
Figure 5. Usefulness rating of the different types of feedback. The median response for each type is shown by the vertical line. The right side of the figure shows pairs of feedback types for which the Durbin-Conover post-hoc test indicates statistically significant differences in ratings. For example, there is a statistically significant difference between the responses for “Automatic testing” and “Score rankings” with $p=1.23\mathrm{e}\text{-}07$.
This study also aims to understand if any type of feedback stood out as particularly more useful than the others. Hence, we asked the students to rank each of the 7 types of feedback from Section 4 from 1 (not useful) to 5 (very useful). Figure 5 presents the answer distribution to these questions. Most students ranked all 7 types of feedback as useful, with all of the medians being $\geq 4$.
Since not all students interacted with all forms of feedback and the Friedman test requires a full block design, we excluded students who did not answer all 7 questions, resulting in 94 samples. The Friedman test allows us to reject the null hypothesis of there being no significant difference between the different types of feedback ($\chi^{2}(6)=58.66,\;p=8.436\times 10^{-11},\;\hat{W}_{Kendall}=0.10$). As such, to discover which types of feedback were significantly more useful, we used the post-hoc pairwise Durbin-Conover test. The right side of Figure 5 shows the pairs of feedback types for which there is a significant difference in median ranks, along with the respective p-values. These statistically significant tests show that “Automatic testing” was the most helpful type of feedback, being more useful than 4 other types, and that “Open choice points warning” and “Predicate scoring” were more useful than 2 other types each. Meanwhile, the least helpful types of feedback were “Score rankings” and “Solution type validation”, with each being less useful than 3 other types.
A1. Most students consider that the provided feedback was helpful and the system has a usability score of B+. The most useful feedback were “Automatic testing”, “Open choice points warning” and “Predicate scoring”, while the least useful were “Score ranking” and “Solution type validation”. These conclusions are supported by statistically significant differences.
6.2. Impact of interest, engagement and LLM use (RQ2)
Another goal of this study is to determine if students’ interest, engagement with optional exercises and having used LLMs (or not) had any effect on their perception of the usefulness of the feedback.
6.2.1. Interest
Figure 6. Answer distribution for the question “How interested are you in the topics of logic and logic programming?”. The median response is shown by the black vertical line.
We asked students to rate their interest in the topics of logic and logic programming on a scale from 1 to 5. The answers, summarized in Figure 6, show a fairly normal distribution, with most students having some interest in the topics and a small number showing no interest or a large amount of interest.
Figure 7. Effect of interest in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
To test our hypothesis that more interested students take better advantage of automated feedback, we decided to compare students’ interest with the agreement level for the statement “I think the automated feedback helped increase my grade in this course”. Figure 7 shows the agreement responses grouped by interest level. A Spearman’s rank correlation analysis does not allow us to reject the null hypothesis, which means there is no statistically significant relationship between students’ interest level and how beneficial they found the feedback received ($S=4.26\times 10^{5},\;p=0.08,\;\hat{\rho}=0.14$).
6.2.2. Engagement
Figure 8. Effect of engagement in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
Students had the opportunity to use ProHelp with two types of optional exercises: a set of 5 introductory exercises and a set of 22 recitation exercises. Considering the 144 respondents to our survey, 47 did not solve any optional exercises, 50 solved at least one exercise from both optional sets, 21 solved only exercises from the introductory set, and 26 solved only exercises from the recitation set. Our hypothesis is that more engaged students (ones who solved optional exercises) would have taken better advantage of ProHelp and thus found it more useful. Figure 8 shows the relationship between the 4 engagement groups and the agreement levels with the statement “I think the automated feedback helped increase my grade in this course”. To test our hypothesis, we used a Kruskal-Wallis test, which shows no evidence of there being a statistical difference between the four groups ($\chi^{2}(3)=2.19,\;p=0.53,\;\hat{\epsilon}^{2}=0.02$).
6.2.3. Usage of LLMs
Table 1. Number of students that reported using each LLM tool.
LLM Tool Count
ChatGPT 82
Copilot 16
Claude 5
Gemini 2
Blackbox AI 1
Deepseek 1
Perplexity AI 1
ZZZ Code AI 1
Figure 9. Effect of LLMs usage in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
Students were allowed to use Large Language Models (LLMs) during this course, as long as they disclosed that usage. Considering the 144 responses to our survey, 91 reported using LLMs when solving assignments, while 53 reported that they did not use LLMs. Table 1 shows the number of students that reported using each LLM tool/platform, with the most popular ones being ChatGPT and Copilot. Furthermore, 21 students reported using more than one LLM. We were interested to know if having used LLMs has an effect on how students perceived the usefulness of ProHelp (for example, students who used LLMs might have found ProHelp less useful due to already having personalized feedback from the LLM). In Figure 9, we show the usefulness ratings for students who used and did not use LLMs. According to the Mann-Whitney $U$ test, there is no statistical difference between the two groups (W=2261.50,p=0.51,r=0.06W=2261.50,\;p=0.51,\;r=-0.06).
A2. There is no evidence that interest level, having used the platform for optional exercises, or having used LLMs, affects how helpful students found ProHelp overall.
6.3. Submission Analysis (RQ3)
Table 2. Number of correct and incorrect submission per assignment.
Assignment Correct Incorrect Total Avg. Clauses
Role-playing Exercise 271 239 510 3.51
    Day 1 88 58 146 3.00
    Day 2 63 67 130 3.27
    Day 3 40 68 108 5.05
    Day 4 46 9 55 2.62
    Day 5 34 37 71 3.26
Recitation Exercises 969 442 1411 2.50
    Exercise A1 65 55 120 2.00
    Exercise A2 79 44 123 2.67
    Exercise A3 61 24 85 2.05
    Exercise A4 49 15 64 2.00
    Exercise A5 49 12 61 3.05
    Exercise A6 50 21 71 2.00
    Exercise A7 46 19 65 2.08
    Exercise B1 57 42 99 3.09
    Exercise B2 39 31 70 1.77
    Exercise B3 42 16 58 3.55
    Exercise B4 41 7 48 1.70
    Exercise B6 45 15 60 2.86
    Exercise C1 41 19 60 2.22
    Exercise C2 36 6 42 2.98
    Exercise C3 35 7 42 3.32
    Exercise C4 37 17 54 3.11
    Exercise D1 47 13 60 2.25
    Exercise D2 35 10 45 2.56
    Exercise D3 36 14 50 3.09
    Exercise D4 33 23 56 3.02
    Exercise D5 23 27 50 2.12
    Exercise D6 23 5 28 2.00
Project 440 4840 5280 34.20
Total 1680 5521 7201
Figure 10. Number of submissions over time for each type of assignment.
Of the 312 students that interacted with ProHelp, 265 gave permission for their code to be used for research purposes. Over the course, those 265 students created a total of 7201 code submissions to the different repositories. We have created a publicly available annotated dataset containing these submissions (Brancas et al., 2025b). Table 2 shows the total number of correct and incorrect submissions per assignment, along with the average number of clauses in the students’ submissions for each one. Overall, there were 1680 submissions that passed all the tests for the respective assignment, and 5521 submissions that failed at least one test. By far, the assignment with the most number of submissions is the final project, which is easily explained by it being the only mandatory and graded assignment. The ratio of correct to incorrect submissions also varies greatly, with the optional assignments having much fewer incorrect submissions than the project. There are two plausible explanations for this: (1) the optional assignments are easier, and thus take fewer tries to get right, and (2) only more engaged students solved the optional exercises. Figure 10 shows the number of submissions for each day in the course and allows us to spot some interesting patterns in students’ behavior. Particularly, that most students only start solving the optional exercises on the platform once the project submissions are opened. It is also possible to see the escalating number of submissions on the three days leading to the project deadline.
Figure 11. Classification of the 7201 student submissions.
We analyzed the 7201 submissions and classified them along two dimensions: (1) is the submission correct or incorrect, and (2) is the submission an improvement or a regression compared with the previous version of the code. For this second category we consider 5 different labels:
  • First Submission: first submission of a student for a particular assignment;
  • Bug Fixed: the submission passes strictly more tests than the previous version;
  • Bug Introduced: the submission passes strictly fewer tests than the previous version;
  • Mixed: the submission both fails tests that previously passed and passes tests that previously failed;
  • No Change: the submission passes exactly the same tests as the previous version.
Figure 11 shows the number of submissions in each category. We observe that most of the first submissions are correct, but, as evidenced by Table 2, this is mostly due to the optional exercises. We also observe that the number of regressions is fairly small (Mixed, Bug Introduced). Lastly, there are a small number of submissions that are purely code quality improvements/refactorings of submissions that were already correct.
In order to understand the types and frequency of bugs committed by students, we analyzed the submissions in the Bug Fixed label. Of those 2495 submissions, we isolated the ones where only one predicate was modified – this subset represents “ideal” unit commits where students fixed a specific incorrect behavior in their program. This selection resulted in 767 submissions, of which we manually classified 200 chosen randomly.
LABEL:tab:bug-types presents the identified bug types and subtypes found in the 200 selected programs, along with their frequencies and a representative example for each type. Note that a single program may contain multiple bugs, and thus the total number of reported bug types exceeds 200. The most commonly identified reason for failing submissions was “Incomplete” programs, which can happen for two reasons: (1) students implemented their programs incrementally and tested components as they progressed (most often represented by the “Missing Predicate” subtype), and (2) students forgot to implement some functionality (most often represented by the “Missing Clause” subtype). While the first case does not necessarily represent a true bug, it still gives us the relevant insight that students seem to like having many tests in order to evaluate each functionality in isolation.
Other bug types were more evenly distributed, with no specific bug occurring at an overwhelmingly high frequency. Of particular note is the “Domain Logic Problem” bug type, which reflects a misunderstanding of the intended functionality for a given predicate. Additionally, 13 of the 26 programs categorized as “Other” contain at least one predicate that is fundamentally incorrect and requires complete rewriting.
A3. We have identified 9 categories of bug types committed by students, with 17 total subtypes. We also created a public dataset with 7201 Prolog programs, including 200 manually annotated with bug types and corrections.
Table 3. Description and frequency of the different bug types identified in student submissions.
Bug Type Count Frequency
Incomplete 75 37.5%
  Subtypes
     Missing Predicate 59 29.5%
     Missing Clause (of existing predicate) 16 8.0%
There are missing predicates/clauses. Example: +mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * N, mult(L1, N, L2).
Wrong Argument 54 27.0%
  Subtypes
     Argument Order Swap 7 3.5%
     Missing Argument 6 3.0%
     Extra Argument 2 1.0%
     Other 39 19.5%
Issue in the arguments of a predicate call. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * N, mult(L1, N, [-L1-]{+L2+}).
Rule Goal Problems 26 13.0%
  Subtypes
     Extra Goal 13 6.5%
     Missing Goal 10 5.0%
     Goal Order Swap 3 1.5%
Problems with the goals in the body of a rule. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * N{+, mult(L1, N, L2)+}.
Operator Error 23 11.5%
  Subtypes
     Wrong Operator 16 8.0%
     List Terminators Issue 4 2.0%
     Missing Negation 3 1.5%
Incorrect operator usage. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 [-=-]{+is+} E1 * N, mult(L1, N, L2).
Wrong Predicate Name 19 9.5%
Incorrect predicate name in call. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * N, mult[-iply-](L1, N, L2).
Domain Logic Problem 16 8.0%
Different problems on the implementation of domain logic.
Cut Problem 14 7.0%
  Subtypes
     Missing Cut 10 5.0%
     Extra Cut 3 1.5%
     Wrong Placement 1 0.5%
Problems with the cut operator (!). Example: max(X, Y, X) :- X >= Y{+, !+}. max(_, Y, Y).
Wrong Variable/Constant 8 4.0%
  Subtypes
     Wrong Variable Name 6 3.0%
     Wrong Constant 2 1.0%
Problems with the names of variables or constants. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * [-M-]{+N+}, mult(L1, N, L2).
Other 26 13.0%
Table 3. (continued)
Figure 12. Usefulness ratings for the different types of feedback not yet implemented in ProHelp. The median response for each type of feedback is shown by the black vertical line. The right side of the figure shows pairs of feedback types for which the Durbin-Conover post-hoc test indicates statistically significant differences in ratings.
6.4. Future directions (RQ4)
Our last goal is to better understand what features and improvements to the feedback students would most like to have in the future, while also analyzing what aspects of the course students did not like. In the survey, we asked students to rate the 4 types of (not yet implemented) feedback described in Section 4.2 between Not Useful, Somewhat Useful and Useful. Figure 12 shows the students’ answers, where at least 75% of students rated each of the proposed feedback types as Useful. These responses hint that these are good goals to work towards.
To check if any type of feedback had a significant preference over another, we used a Friedman test which rejects the null hypothesis ($\chi^{2}(3)=20.74,\;p=1.19\times 10^{-4},\;\hat{W}_{Kendall}=0.05$). Next, to check which types of feedback students think will be more useful, we used the post-hoc pairwise Durbin-Conover test. The right side of Figure 12 shows the pairs of feedback types that showed statistical significance. Students rated “Showing the differences between generated and expected output” as significantly more useful than the other 3 types of feedback.
Table 4. Description of issues and features raised by students in open-ended responses, along with the count of how many students voiced each one. Issues/features reported by only one student were omitted.
Description Count
Predicate scores should be proportional to the number of tests passed 11
Tests should be more specific and in greater quantity 10
The project statement and/or organization of the repository was confusing 5
It should be clear what predicates each test is targeting 4
Having more feedback might not lead to better learning 3
Instructions for how to use Git/VS Code/etc were confusing 3
The faculty was not very responsive on Discord 3
Feature: show test execution times 2
Feature: predict the qualitative grade portion of the evaluation 2
Waiting 15 minutes between submissions is too long 2
The way submissions are counted was unfair 2
At the end of the survey, we asked students to suggest other types of feedback they would like to see in the future, as well as provide general comments/criticisms. Table 4 shows the issues raised by students. One common criticism among students was that the unit tests were not specific enough, i.e., each test covered too much functionality. For example, P-31 said: “The project tests were not exhaustive enough.”. This left students with little information when tests failed and highlights the importance of faculty creating good unit tests that assess each functionality in isolation. There were also other comments regarding the course organization and grading method, which we have relayed to the faculty.
One interesting suggestion provided by some students regards the qualitative component of the project grade, which is manually attributed and encompasses characteristics such as code readability, organization and documentation. Students suggest that it would be useful to have an estimate for this qualitative component in the automated feedback, even if the final score continued to be hand-assigned by faculty.
A4. Although students found all proposed types of feedback useful, there is a significant preference for the “Show the outputs and expected outputs” compared with the remaining feedback types. Students would like to have an automated estimation for the qualitative component of their projects.
7. Discussion and Conclusion
This work analyses students’ behaviors during a 9-week long undergraduate class on logic programming. Specifically, we study how they interacted with an automated assessment platform for Prolog code. This platform, named ProHelp, included several types of automated feedback to guide students to find and solve their bugs more easily. Our main contributions are: the analysis of a student survey with 144 valid responses, gathering insights about students’ preferences and complaints regarding ProHelp, and a dataset of 7201 Prolog programs with 200 manually annotated ones with bug types and corrections.
Survey analysis indicates that users found the feedback received useful and the ProHelp interface was user-friendly. For instance, P-101 said “I loved the assessment system. It was very efficient and helpful. From what I have heard from previous years, it was a huge improvement”. The components of ProHelp that students most appreciated include automatic testing, warnings about open choice points, and scoring for each required predicate. Students also displayed great interest in the more advanced types of feedback we proposed and suggested some themselves. These findings will be valuable for future research, enabling educators and developers to focus on the most impactful types of feedback.
Results show no statistically significant relationship between students’ perception of feedback usefulness and their interest, engagement, or usage of LLMs. Even so, there is a small but non-statistically significant correlation ($p=0.08$) between students’ interest and their usefulness rating, suggesting that further research – with a larger sample or different methodology – is warranted. The findings on LLM usage are particularly intriguing, and we propose several possible explanations for it: (1) students under-reported LLM-usage due to its negative connotation, (2) the feedback received in ProHelp is orthogonal to the questions students usually ask LLMs, or (3) LLMs are not (yet) very useful for Prolog. Further investigation is needed to determine which, if any, of these factors apply.
Our submission code analysis resulted in a public dataset which can be used by future work to develop and test fault localization and program repair techniques. The report on the types of bugs students commit most frequently can also be used by faculty in order to stress those point a priori in-class. Likewise, insights such as students only starting to solve optional exercises once that knowledge is also required for graded work can also be used by faculty in the development of future courses.
Building on the results of the survey and the submissions dataset, we plan to enhance ProHelp with the following improvements: (1) use fault localization and program diagnosis techniques to give more information to students when tests fail, particularly for failed assertions and infinite loops, (2) explore automated repair methods, namely program synthesis and LLMs, to generate correction hints or bug explanations for students, and (3) investigate LLM-based techniques to estimate qualitative aspects of logic programs. One challenge highlighted by some students is that excessive feedback may hinder learning rather than help. Balancing the amount and type of feedback will be crucial to ensuring that students receive meaningful guidance without becoming overly reliant on automation.
Acknowledgements.
This work was partially supported by Portuguese national funds through FCT, under project 2023.14280.PEX (DOI: 10.54499/2023.14280.PEX), and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (FUN2MODEL, grant agreement No. 834115). This work was also supported by the Carnegie Mellon University Portugal Program and FCT under grants PRT/BD/152086/2021 (DOI: 10.54499/PRT/BD/152086/2021) and PRT/BD/153739/2021 (DOI: 10.54499/PRT/BD/153739/2021).
Appendix A Survey Protocol
The following survey was designed to collect anonymous feedback from undegraduate students enrolled in the Logic Programming course regarding their experience with the automated feedback system provided via GitLab. The survey was estimated to take approximately 10 to 15 minutes. Several questions were included to assess usability, usefulness, and pedagogical effectiveness of the feedback tools. Some questions functioned as attention checks. The questions and answers presented here were translated from Portuguese.
Section 1: Demographics and Background
  • What gender do you identify with?
    • Female
    • Male
    • Prefer not to say
    • Prefer to self-describe
  • What is your level of interest in the topics covered in the Programming Languages course? (Logic, Prolog)
    Rating scale from 1 (not interested) to 5 (very interested)
  • During this course, did you experience difficulties using Git/GitLab? Select all that apply:
    • Merge conflicts
    • SSH key problems
    • Problems interacting with VS Code
    • I had no problems
    • Other
Section 2: Automated Feedback in Prolog
Students were reminded of the types of automated feedback they encountered on GitLab:
  • Automatic test verification
  • Predicate scoring
  • Validation of solution types
  • Dashboards with scores
  • Syntax error highlighting
  • Suggestions for incorrect predicate names
  • Warnings about open choice points
Students were asked to rate their agreement with the following System Usability Survey Items using a 5-point Likert scale (1 = Strongly Disagree, 5 = Strongly Agree):
  • I think that I would like to use this system frequently.
  • I found the system unnecessarily complex.
  • I thought the system was easy to use.
  • I think that I would need the support of a technical person to be able to use this system.
  • I found the various functions in this system were well integrated.
  • I thought there was too much inconsistency in this system.
  • I would imagine that most people would learn to use this system very quickly.
  • I found the system very cumbersome to use.
  • I felt very confident using the system.
  • I needed to learn a lot of things before I could get going with this system.
Additionally, students rated the usefulness of each feedback component on a scale from 1 (Not useful at all) to 5 (Very useful):
  • Suggestions for incorrect predicate names
  • Syntax error highlighting
  • Dashboards with scores
  • Select “Not useful at all” in this line [attention check]
  • Automatic test verification
  • Predicate scoring
  • Warning for open choice points
  • Solution type validation (iterative, recursive, functional)
Section 4: Role-playing Exercise
  • Did you solve the role-playing exercise?
    • No
    • I started but did not finish
    • Yes
  • Why did you not solve/abandon/finish the exercises? (Open text response)
  • Did you use LLMs (e.g., ChatGPT, Copilot) to help solve the role-playing exercise?
    • No
    • Yes
  • Which LLM(s) did you use? (Open text response)
  • What did you think of the sequence of puzzles from the role-playing exercise on a 5-point scale (1 = Not entertaining at all, 5 = Very entertaining)?
Section 5: Recitation Exercises
  • Did you solve the recitation exercises using Git/GitLab?
    • No
    • I started but did not finish
    • Yes
  • Why did you not solve/abandon/finish the exercises? (Open text response)
  • Did you use LLMs (e.g., ChatGPT, Copilot) to help with the recitation exercises?
    • No
    • Yes
  • Which LLM(s) did you use? (Open text response)
Section 6: Project
  • Did you use LLMs (e.g., ChatGPT, Copilot) to help with the project?
    • No
    • Yes
  • Which LLM(s) did you use? (Open text response)
Section 7: Final Reflections
  • Please rate the following statements on a 5-point Likert scale (1 = Strongly Disagree, 5 = Strongly Agree)
    • I believe the automated feedback I received on GitLab improved my grade in this course.
    • I believe the role-playing exercise improved my grade in this course.
    • Select “Strongly Agree” to demonstrate you are paying attention. [Attention check]
    • I believe the recitation exercises on GitLab improved my grade in this course.
  • Please rate the potential utility of proposed features on a 3-point scale (Indifferent, Somewhat Useful, Useful):
    • Suggestions for corrections on incorrect submissions
    • Showing program output for each test compared to expected output
    • Automatic identification of bug locations
    • Detection of infinite loops
  • Please suggest any other features you would like to see. (Open text response)
  • Please leave any other feedback you may have (positive or negative). (Open text response)